# The expected value of sample information calculations for external validation of risk prediction models

### v2

@@ Methods
The power of this approach is in the flexibility of the bootstrap method in accommodating different types of outcomes and practical considerations in validation studies. For example, if the dataset $\mathbf{d}$ has non-trivial level of missing data, which affects the amount of information it contains, the step 1b in the above algorithm can include an imputation component, generating the population $\mathbf{F}$ without any missing values (and similarly, step 1d can introduce missingness in the data to simulate a real-world validation dataset).
@@
The ordinary bootstrap, which assigns weights from a scaled $\text{Multinomial}(1/n,1/n,...)$ distribution to observations, is conceptually similar to the Bayesian bootstrap and has been interpreted in a Bayesian way (e.g., in the approximate Bayesian bootstrap employed in missing value imputation23). In data-driven VoI analyses it generates similar results to the Bayesian bootstrap11,22. It thus can be used in lieu of the Bayesian bootstrap in step 1a.
@@
An algorithm based on beta-binomial conjugacy for binary outcomes
@@
In their proposal for a parametric Bayesian NB estimation for binary outcomes, Cruz and Korthauer take advantage of the beta-binomial conjugacy and the fact that the likelihood function for $NB$ can be factorized into terms involving prevalence, sensitivity, and specificity. As such, they argue that specifying prior information in terms of independent beta distributions on these quantities will result in corresponding posterior distributions that are also independent beta20. In this case, given that $P(\theta|\mathbf{d})$ is composed of three independent beta distributions, by the same token $P(\theta|\mathbf{d},\mathcal{D^{*}})$ after a realization of $\mathcal{D^{*}}$ will also be the product of three independent beta distributions. This results in a fast Monte-Carlo approach for EVSI computation as explained in Table 2.
0. Specify the current information for NBs under current information as $\theta_{p}\sim\text{Beta}(\alpha_{p},\beta_{p})$, $\theta_{se}\sim\text{Beta}(\alpha_{se},\beta_{se})$, $\theta_{sp}\sim\text{Beta}(\alpha_{sp},\beta_{sp})$. 1. Plug in the expected values of $\theta$ from the above distributions in equation 1. Record the maximum expected NB. This is $\mathbb{E}NB_{\text{current information}}$. 2. For $j$ = 1 to $N$ (number of Monte Carlo simulations) (a) Randomly draw from the distribution of $\theta$ as specified in step 2. (b) Compute true NBs from such draws. Record the maximum NB. (c) Generate $\mathcal{D^{*}}$, defined by $\{n^{*}_{tp},n^{*}_{fn},n^{*}_{tn},n^{*}_{fp}\}$ given the draw from $\theta$ obtained in step 2a as: $n^{*}_{+}\sim\text{Binomial}(n^{*},\theta_{p})$ (number of positive cases in the future sample), $n^{*}_{tp}\sim\text{Binomial}(n^{*}_{+},\theta_{se})$, $n^{*}_{fn}=n^{*}_{+}-n^{*}_{tp}$, $n^{*}_{tn}\sim\text{Binomial}(n^{*}-n^{*}_{+},\theta_{sp})$, $n^{*}_{fp}=n^{*}-n^{*}_{+}-n^{*}_{tn}$. (d) Calculate expected NBs given generated $\mathcal{D^{*}}$ by plugging in the updated estimates of the expected value of $\theta$ as follows: ${\bar{\theta}^{*}_{p}}=(\alpha_{p}+n^{*}_{tp}+n^{*}_{fn})/(\alpha_{p}+\beta_{p}+n^{*})$, ${\bar{\theta}^{*}_{se}}=(\alpha_{se}+n^{*}_{tp})/(\alpha_{se}+\beta_{se}+n^{*}_{tp}+n^{*}_{fn})$, ${\bar{\theta}^{*}_{sp}}=(\alpha_{sp}+n^{*}_{tn})/(\alpha_{sp}+\beta_{sp}+n^{*}_{tn}+n^{*}_{fp})$. Record the maximum NB. This will be $\mathbb{E}_{\mathbf{\theta|\mathbf{d},\mathcal{D^{*}}}}NB(i,\mathbf{\theta})$. Next $j$ 3. Average NBs from step 2b. This is $\mathbb{E}NB_{\text{perfect information}}$. From this subtract ENB from step 1. This is EVPI. 4. Average NBs from step 2d. This is $\mathbb{E}NB_{\text{sample information}}$. From this subtract ENB from step 1. This is EVSI.
Table 2: Computation of EVSI based on beta-binomial conjugacy
@@
Independent beta distributions for each component of $\theta$ can emerge, from example, if we ask an expert to express their belief about prevalence, sensitivity, and specificity, in terms of fractions. That is, if their best estimate of prevalence is 30%, and their uncertainty is akin to their opinion comping from a sample of 10 individuals, the information can be expressed as $\theta_{p}\sim\text{Beta}(3,7)$ (similar for sensitivity and specificity).
@@
Importantly, for binary outcomes and in the absence of missing data, the Bayesian bootstrap approach explained above reduces to this beta-binomial model. This is because the information in a sample $\mathbf{d}$ can be fully specified via four quantities: the number of true positives ($n_{tp}$), false negatives ($n_{fn}$), true negatives ($n_{tn}$), and false positives ($n_{fp})$. The Bayesian bootstrapping of individual observations is akin to assigning an improper prior $\text{Dirichlet}(0,0,0,0)$ to these quantities, giving rise to the posterior distribution being of the form $\text{Dirichlet}(n_{tp},n_{fn},n_{tn},n_{fp})$. Due to the aggregate property of the Dirichlet distribution24, the $\text{Dirichlet}(0,0,0,0)$ prior of the Bayesian bootstrap is equal to assigning $\text{Beta}(0,0)$ priors to $\theta_{p}$, $\theta_{se}$, and $\theta_{sp}$. As such, the beta-binomial algorithms will be equal to the Bayesian bootstrap and can be used instead for VoI computations due to its computational efficiency.
@@
Yet another utility of this approach is that instead of the improper prior employed in the Bayesian bootstrap, the user can specify informative priors. For example, if an independent cross-sectional study in the same target population has estimated outcome prevalence that can be expressed as $\text{Beta}(\alpha,\beta)$, then combined with observed data, our current information about prevalence can be expressed as $\theta_{p}\sim\text{Beta}(\alpha+n_{tp}+n_{fn},\beta+n_{fp}+n_{tn})$. Further, even in the absence of prior information, proper non-informative or weakly informative priors can be used to rectify some of the previously-identified problems with bootstrap-based VoI calculations at extreme situations. In small validation samples and extreme thresholds, VoI calculations might generate counter-intuitive results11. One instance is when sample estimates of $\theta_{se}$ or $\theta_{sp}$ are at their extreme (0 or 1). The improper $\text{Beta}(0,0)$ prior on sensitivity and specificity can result in improper beta posteriors with one of their two parameters being 0. Therefore, we suggest specifying proper priors (e.g., a flat prior $\text{Beta}(1,1)$) for the three quantities in $\theta$ when current information is from an observed sample.
@@
In many practical situations our current information $P(\theta|\mathbf{d})$ may not have a simple mathematical form, but one can still draw a Monte Carlo sample of arbitrary size from its distribution (e.g., via Markov Chain Monte Carlo methods). Examples include when $P(\theta_{p})$ is constructed from a meta-analysis of prevalence studies, or $P(\theta_{se},\theta_{sp})$ from a bivariate meta-analysis of the diagnostic accuracy of the test, giving rise to a joint distribution for $P(\theta_{se},\theta_{sp})$, potentially conditional on prevalence25.
@@
For such a general case where we have a sample from $P(\theta|\mathbf{d})$, the empirical distribution of the sample can be used as the distribution for the current information. The posterior distribution after observing future data can then be approximated by the discrete distribution that puts a mass $w^{(i)}\propto P(\mathcal{D}^{*}|\theta^{(i)})$ (e.g., a product of Binomial probabilities for binary outcomes) on each observation from this sample. This will result in the generic algorithm explained in Table 3.
0. Quantify current information given arbitrary distribution $P(\theta|\mathbf{d})$. Obtain $M$ draws $(\theta^{(1)},\theta^{(2)},...,\theta^{(M)})$ from $P(\theta|\mathbf{d})$. 1. Estimate expected NBs under current information by taking expectation of NB terms (in equation 1) with respect to the empirical distribution that puts a mass $w^{(k)}$ of $\propto 1$ on each $\theta^{(k)}$: $NB(i,\mathbf{\theta})=\left\{\begin{array}[]{ll}0&i=0\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}\theta^{(k)}_{se}-(1-\theta^{(k)}_{p})(1-\theta^{(k)}_{sp})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=1\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}-(1-\theta^{(k)}_{p})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=2\\ \end{array}\right..$ Record the maximum expected $NB$. This is $\mathbb{E}NB_{\text{current information}}$. 2. For $j$ = 1 to $N$ (number of Monte Carlo simulations) (a) Obtain a random draw from $P(\theta|\mathbf{d})$ (or loop over the draws from step 3). This is taken as the true $\theta$ in this iteration. (b) Calculate true NBs given $\theta$ obtained in the previous step. Record the maximum NB. (c) Generate $\mathcal{D^{*}}$ of size $n^{*}$, defined by $\{n^{*}_{tp},n^{*}_{fn},n^{*}_{tn},n^{*}_{fp}\}$ given the sample from $\theta$ obtained in step 2a: $n^{*}_{+}\sim\text{Binomial}(n^{*},\theta_{p})$ (number of positive cases in the future sample), $n^{*}_{tp}\sim\text{Binomial}(n^{*}_{+},\theta_{se})$, $n^{*}_{fn}=n^{*}_{+}-n^{*}_{tp}$, $n^{*}_{tn}\sim\text{Binomial}(n^{*}-n^{*}_{+},\theta_{sp})$, $n^{*}_{fp}=n^{*}-n^{*}_{+}-n^{*}_{tn}$. (d) Update $P(\theta|\mathbf{d})$ by updating the weights $w^{(k)}$ in the empirical distribution of $\theta|\mathbf{d}$ (from step 1) based on the likelihood $P(\mathcal{D^{*}}|\theta)$. For binary outcomes the weights are: $w^{(k)}\propto P_{Bin}(n^{*}_{+},n^{*},\theta_{p}^{(k)})\times P_{Bin}(n^{*}_{tp},n^{*}_{+},\theta_{se}^{(k)})\times P_{Bin}(n^{*}_{tn},n^{*}-n^{*}_{+},\theta_{sp}^{(k)})$, with $P_{Bin}(x,m,p)$ being the probability mass function of the binomial distribution at value $x$ for $m$ trials and success probability $p$. (e) Calculate updated NBs given future data by taking expectation of NB terms in step 1 using the updated weights from step 2d. Next $j$ 3. Average NBs from from step 2b. This is $\mathbb{E}NB_{\text{perfect information}}$. Subtract from this ENB from step 1. This is EVPI. 4. Average NBs from step 2e. This is $\mathbb{E}NB_{\text{sample information}}$. Subtract from this ENB from step 1. This is EVSI.

### v3

@@ 1. For $j$ = 1 to $M$ (number of Monte Carlo simulations)
The power of this approach is in the flexibility of the bootstrap method in accommodating different types of outcomes and practical considerations in validation studies. For example, if $\mathbf{d}$ has non-trivial level of missing data, which affects the amount of information it contains, Step 1c in Table 1 can include an imputation component, generating the population $\mathcal{D^{*}}$ without any missing values (and similarly, Step 1d can introduce missingness in the future sample to simulate a real-world validation dataset).
@@
The ordinary bootstrap, which assigns weights from a scaled $\text{Multinomial}(1,1,...,1)$ distribution to observations in $\mathbf{d}$, is conceptually similar to the Bayesian bootstrap and has been interpreted in a Bayesian way (e.g., in the approximate Bayesian bootstrap employed in missing value imputation27). In data-driven VoI analyses it generates similar results to the Bayesian bootstrap25. It thus can be used in lieu of the Bayesian bootstrap in step 1a.
@@
A fast algorithm based on beta-binomial modeling for binary outcomes
@@
In their proposal for a parametric Bayesian NB estimation for binary outcomes, Cruz and Korthauer note that the likelihood function for $NB$ can be factorized into terms involving prevalence, sensitivity, and specificity28. As such, they take advantage of the beta-binomial conjugacy and note that specifying prior information in terms of independent beta distributions on these quantities will result in corresponding posterior distributions that are also independent beta. In this case, given that $P(\theta|\mathbf{d})$ is composed of three independent beta distributions, by the same token $P(\theta|\mathbf{d},\mathcal{D})$ after a realization of $\mathcal{D}$ will also be composed of three independent beta distributions. This results in a fast Monte Carlo approach for EVSI computation as explained in Table 2.
- Prerequisite: Specify current information $P(\theta|\mathbf{d})$ as
$\theta_{p}\sim\text{Beta}(\alpha_{p},\beta_{p})$, $\theta_{se}\sim\text{Beta}(\alpha_{se},\beta_{se})$, $\theta_{sp}\sim\text{Beta}(\alpha_{sp},\beta_{sp})$.
1. For $j$ = 1 to $M$ (number of Monte Carlo simulations)
(a) Obtain $\theta^{*}$, by random drawing from $P(\theta|\mathbf{d})$. These are true $\theta$s in this iteration.
(b) Calculate $NB^{*}_{0}$, $NB^{*}_{1}$, and $NB^{*}_{2}$, by plugging $\theta^{*}$ into Equation 1. These are true NBs in this iteration. Record the maximum true NB: $NB^{*}_{\text{truth}}=\max\{NB^{*}_{0},NB^{*}_{1},NB^{*}_{2}\}$.
(c) Generate $\mathcal{D}$, the sample of future study, defined by $\{\mathit{N}_{tp},\mathit{N}_{fn},\mathit{N}_{tn},\mathit{N}_{fp}\}$ given $\theta^{*}$ obtained in Step 1a as:
$\mathit{N}_{+}\sim\text{Binomial}(N,\theta_{p}^{*})$ (number of positive cases in the future sample),
$\mathit{N}_{tp}\sim\text{Binomial}(\mathit{N}_{+},\theta_{se}^{*})$,
$\mathit{N}_{fn}=\mathit{N}_{+}-\mathit{N}_{tp}$,
$\mathit{N}_{tn}\sim\text{Binomial}(\mathit{N}-\mathit{N}_{+},\theta_{sp}^{*})$,
$\mathit{N}_{fp}=\mathit{N}-\mathit{N}_{+}-\mathit{N}_{tn}$.
(d) Calculate $\theta^{+}$, the updated estimates of prevalence, sensitivity, and specificity after obtaining the future sample. :
${\theta^{+}_{p}}=(\alpha_{p}+\mathit{N}_{tp}+\mathit{N}_{fn})/(\alpha_{p}+\beta_{p}+\mathit{N})$,
${\theta^{+}_{se}}=(\alpha_{se}+\mathit{N}_{tp})/(\alpha_{se}+\beta_{se}+\mathit{N}_{tp}+\mathit{N}_{fn})$,
${\theta^{+}_{sp}}=(\alpha_{sp}+\mathit{N}_{tn})/(\alpha_{sp}+\beta_{sp}+\mathit{N}_{tn}+\mathit{N}_{fp})$.
(e) Calculate $NB^{+}_{0}$, $NB^{+}_{1}$, and $NB^{+}_{2}$, updated estimates of NBs after obtaining the future sample, by plugging $\theta^{+}$ into Equation 1. Determine the strategy that has the highest expected future NB: $l={\mbox{argmax}_{i\in\{0,1,2\}}(NB^{+}_{i})}$. Record the true NB of this strategy: $NB^{*}_{\text{sample}}=NB^{*}_{l}$.
- Next $j$
2. Take the average of $NB^{*}$s from step 1b and pick the maximum value. This is ${{E}}NB_{\text{current}}$.
3. Average $NB^{*}_{\text{truth}}$s from step 1b. This is ${{E}}NB_{\text{truth}}$. From this subtract ${{E}}NB_{\text{current}}$. This is EVPI.
4. Average $NB^{*}_{\text{sample}}$s from step 1e. This is ${{E}}NB_{\text{sample}}$. From this subtract ${{E}}NB_{\text{current}}$. This is EVSI.
Table 2: Computation of EVSI based on beta-binomial modeling
@@
Independent beta distributions for each component of $\theta$ can emerge, for example, if we ask an expert to express their belief about prevalence, sensitivity, and specificity, in terms of fractions. That is, if their best estimate of prevalence is 30%, and their uncertainty is akin to their opinion comping from a sample of 10 individuals, the information can be expressed as $\theta_{p}\sim\text{Beta}(3,7)$ (similar for sensitivity and specificity).
@@
Importantly, for binary outcomes and in the absence of missing data, the Bayesian bootstrap approach explained previously reduces to this beta-binomial model. This is because the information in a sample $\mathbf{d}$ for estimating NBs can be fully specified via four quantities: the number of true positives ($n_{tp}$), false negatives ($n_{fn}$), true negatives ($n_{tn}$), and false positives ($n_{fp})$. Due to the aggregate property of the Dirichlet distribution29, the $\text{Dirichlet}(0,0,...,0)$ prior used in the Bayesian bootstrap is equal to assigning $\text{Beta}(0,0)$ priors to $\theta_{p}$, $\theta_{se}$, and $\theta_{sp}$, resulting in the posterior distribution for the population that can be specified as $(p_{tp},p_{fn},p_{tn},p_{fp})\sim\text{Dirichlet}(n_{tp},n_{fn},n_{tn},n_{fp})$. As such, the beta-binomial algorithm will be equal to the Bayesian bootstrap and can be used instead for VoI computations due to its computational efficiency.
@@
Yet another utility of this approach is that instead of the improper prior employed in the Bayesian bootstrap, the user can specify informative priors. Imagine in the current sample $\mathbf{d}$ of size $n$, there are $m$ individuals who have experienced the outcome. If an independent cross-sectional study in the same target population has estimated outcome prevalence that can be expressed as $\text{Beta}(\alpha,\beta)$, our current information about prevalence can be expressed as $\theta_{p}\sim\text{Beta}(\alpha+m,\beta+n-m)$. Even in the absence of external information, weakly informative priors can be employed to rectify some of the previously-identified problems with bootstrap-based VoI calculations at extreme situations. In small validation samples and extreme thresholds, VoI calculations might generate counter-intuitive results15. One instance is when sample estimates of $\theta_{se}$ or $\theta_{sp}$ are at their extreme (0 or 1). The improper $\text{Beta}(0,0)$ prior on these quantities results in improper beta posteriors with one of their two parameters being 0. This can be resolved if one uses a flat $\text{Beta}(1,1)$ prior for these quantities.
@@
In many practical situations our current information $P(\theta|\mathbf{d})$ may not have a simple mathematical form, but one can still draw a Monte Carlo sample of arbitrary size from its distribution. Examples include when $P(\theta_{p})$ is constructed from a meta-analysis of prevalence studies, or $P(\theta_{se},\theta_{sp})$ from a bivariate meta-analysis of the diagnostic accuracy of the test, giving rise to a joint distribution for $P(\theta_{se},\theta_{sp})$30. Markov-Chain Monte-Carlo methods can be used to obtain samples from the joint distribution of $\theta$, even though the mathematical form of the joint density might be intractable (an example is the case studies in Wynants et al21).
@@
For such a general case where we have a sample from $P(\theta|\mathbf{d})$, the empirical distribution of the sample can be used as the distribution for the current information. We note that $P(\theta|\mathbf{d},\mathcal{D})\propto P(\theta)P(\mathbf{d},\mathcal{D}|\theta)=P(\theta)P(\mathbf{d}|\theta)P(\mathcal{D}|\theta)\propto P(\theta|\mathbf{d})P(\mathcal{D}|\theta)$. As such, the posterior distribution of $\theta$ after observing the future sample can be approximated by the discrete distribution that puts a mass $w^{(i)}\propto P(\mathcal{D}|\theta^{(i)})$ on the $i$th observation of the sample from $P(\theta|\mathbf{d})$. This will result in the general algorithm explained in Table 3.
- Prerequisite: Specify current information as arbitrary distribution $P(\theta|\mathbf{d})$. Obtain $M$ draws $(\theta^{(1)},\theta^{(2)},...,\theta^{(M)})$ from $P(\theta|\mathbf{d})$.
1. For $j$ = 1 to $M$ (number of Monte Carlo simulations)
(a) Let $\theta^{*}=\theta^{(i)}$. This is true $\theta$ in this $\dagger$.
(b) Calculate $NB^{*}_{0}$, $NB^{*}_{1}$, and $NB^{*}_{2}$, by plugging $\theta^{*}$ into Equation 1. These are true NBs in this iteration. Record the maximum true NB: $NB^{*}_{\text{truth}}=\max\{NB^{*}_{0},NB^{*}_{1},NB^{*}_{2}\}$.
(c) Generate $\mathcal{D}$, the sample of future study, defined by $\{\mathit{N}_{tp},\mathit{N}_{fn},\mathit{N}_{tn},\mathit{N}_{fp}\}$ as follows:
$\mathit{N}_{+}\sim\text{Binomial}(\mathit{N},\theta^{*}_{p})$ (number of positive cases in the future sample),
$\mathit{N}_{tp}\sim\text{Binomial}(\mathit{N}_{+},\theta^{*}_{se})$,
$\mathit{N}_{fn}=\mathit{N}_{+}-\mathit{N}_{tp}$,
$\mathit{N}_{tn}\sim\text{Binomial}(\mathit{N}-\mathit{N}_{+},\theta^{*}_{sp})$,
$\mathit{N}_{fp}=\mathit{N}-\mathit{N}_{+}-\mathit{N}_{tn}$.
(d) Update the distribution of $\theta$ given the future sample. This is done by updating the weights $w^{(k)}$ based on the likelihood $P(\mathcal{D}|\theta)$. For binary outcomes the weights are:
$w^{(k)}\propto P_{Bin}(\mathit{N}_{tp}+\mathit{N}_{fn},\mathit{N},\theta_{p}^{(k)})\times P_{Bin}(\mathit{N}_{tp},\mathit{N}_{tp}+\mathit{N}_{fn},\theta_{se}^{(k)})\times P_{Bin}(\mathit{N}_{tn},\mathit{N}_{tn}+\mathit{N}_{fp},\theta_{sp}^{(k)})$,
with $P_{Bin}(x,m,p)$ being the probability mass function of the binomial distribution at value $x$ for $m$ trials and success probability $p$.
(e) Calculate $NB^{+}_{0}$, $NB^{+}_{1}$, and $NB^{+}_{2}$, updated estimates of expected NBs after obtaining the future sample, by using the weights generated in the previous step as follows:
$NB^{+}_{i}=\left\{\begin{array}[]{ll}0&i=0\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}\theta^{(k)}_{se}-(1-\theta^{(k)}_{p})(1-\theta^{(k)}_{sp})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=1\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}-(1-\theta^{(k)}_{p})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=2\\ \end{array}\right..$
Determine the strategy that has the highest expected NB in this pooled sample: $l={\mbox{argmax}_{i\in\{0,1,2\}}(NB^{+}_{i})}$. Record the true NB of this strategy: $NB^{*}_{\text{sample}}=NB^{*}_{l}$.
- Next $j$
2. Average $NB^{*}_{\text{truth}}$s from from step 1b. This is ${{E}}NB_{\text{truth}}$. Subtract from this ${{E}}NB_{\text{current}}$. This is EVPI.
3. Average $NB^{*}_{\text{sample}}$s from step 1e. This is ${{E}}NB_{\text{sample}}$. Subtract from this ${{E}}NB_{\text{current}}$. This is EVSI.
$\dagger$ If the size of the sample is large it might not be computationally feasible to loop over all observations. In which case one can pick one of $\theta^{(i)}s$ at random in this step, but still uses the entire sample in step 1e.
@@
Case Study
We used data from GUSTO-I, a clinical trial of multiple thrombolytic strategies for acute myocardial infarction (AMI), as our case study31. This dataset has been widely used for methodological research in predictive analytics32–34. All analyses are conducted in R35, with a fast implementation of the general EVSI algorithm in C++ (please refer to the evsiexval R package for details).
For the present case study, we use a similar setup as our previous work, with a risk prediction model for 30-day mortality developed using the non-US sample of this dataset, and validated in the US sub-sample. The risk prediction model is a logistic regression model fitted to the entire non-US subset (sample size 17,796). We initially assume we have access to only $n=500$ individuals from the US sub-sample as the source of current information. We randomly selected 500 individuals, without replacement, from the entire 23,034 observations in the US sub-sample of GUSTO-I. Later, we will use more observations from the US sub-sample to study how EVSI behaves with the amount of current information. We consider the range of relevant thresholds to be 1%-10%, with 1%, 2%, 5%, and 10% being of primary interest. VoI computations are based on 10^6 Monte Carlo simulations using the aggregate beta-binomial model, with a flat $\text{Beta}(1,1)$ prior on each component of $\theta$.
Using this setup, we performed two simulation studies to explore the face validity of the proposed algorithms. In the first set, we aimed at comparing the numerical stability and computational time of the three algorithms. In the second set, we explored how EVSI changes as a function of the amount of current information, represented by $n$, the size of the current validation data from which $P(\theta|\mathbf{d})$ is constructed.